# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v1

@@ 5 Complexity Analysis of Foreplan
Note that $c$ is bounded by the number of PRVs in our parfactor model, as one parfactor is sufficient per PRV. Theorem 12 overapproximates the size of the state representation, as not all cliques have the same size and not both, $c$ and $w$ are large at the same time. Also, $c$ and $w$ are determined by the structure of the relational cost graph and independent of the domain sizes. Building on the size of the state representation, we give the complexity of the state and action space:
Theorem 13.
The state and action spaces are both polynomial in the number of objects and exponential in $c$ and $w$.
@@
We need to iterate over all possible instantiations of the state representation. For each clique (resp. CRV), the number of possible instantiations is polynomial in the number of objects. The joint state requires one instantiation per clique, resulting in the number of PRVs as exponent. The size of the action space is bounded by the size of the state space, as the action has to specify a (subset of a) state. ∎
Since Foreplan uses a linear program to compute the value function, we analyze the complexity of solving the linear program Foreplan builds. Linear programs can be solved in polynomial time w.r.t the variables and constraint [33]. Let us therefore take a closer look at the number of constraints and variables Foreplan generates:
Theorem 14.
@@
Plugging Theorem 13 in Theorem 14 leads to:
Theorem 15.
The runtime of Foreplan is polynomial in the number of objects and exponential in $c$ and $w$.
For lifting, the runtime in the number of objects is of most interest, because all other model parameters are mostly fixed. Therefore, we have successfully developed a lifted forward planner in rfMDPs. Foreplan is an exact algorithm and efficient w.r.t domain sizes. But, if an approximate solution is good enough and speed is of great concern, we can use an approximate version of Foreplan which is blazingly fast. In the next section, we introduce the approximate version, using approximation to prevent iterating the whole state space and thus circumventing the exponential influence of $c$.
6 Foreplan: Faster by Approximation
While Foreplan runs in time polynomial in the number of objects, the runtime still depends exponentially on $c$. In this section, we present an approximation technique inspired by the Approximate Linear Programming (ALP) [16] approach to prevent iterating the whole state space. We first describe the approximation idea and then how Foreplan uses the approximation. Last, we give bounds on the runtime and on the approximation quality. We call the approximate version Approximate Foreplan.
Foreplan needs to iterate over the whole state space, because the value function maps from a state to the value of that state. We approximate the value function by a set $h_{i}$ of basis functions, whose scope is a subset of $\mathbf{S}$, $V\approx\sum_{i}w_{i}h_{i}$, where the goal is to find the most suitable weights $w_{i}$ [16]. Approximate Foreplan also needs the value of all possible next states in terms of the same approximation. Therefore, Approximate Foreplan makes use of the backprojections $g_{i}^{a}$ of the basis functions $h_{i}$ [16], stating the influence of $x$ on the next state:
@@ 6 Foreplan: Faster by Approximation
Approximate Foreplan parameterizes the basis functions like the rewards to compute the basis functions and backprojections lifted:
Example 10 (Basis Function).
We have three basis functions: $h_{0}=1$, $h_{1}(Sick(M))\coloneq R_{1}(Sick(M))$ and $h_{2}(Travel(M))\coloneq R_{2}(Travel(M))$.
The basis functions should capture the important dynamics in the model [21]. The backprojections are computed lifted as defined next:
Definition 16 (Lifted Backprojection).
Given a basis function $h_{i}$ and Boolean assignments $\tilde{x}$ and $\tilde{a}$ to the state and action, respectively, the backprojection is defined as $g_{i}^{\tilde{a}}(\tilde{x})=\sum_{\tilde{x}^{\prime}}P(\tilde{x}^{\prime}\mid\tilde{x},\tilde{a})\cdot h_{i}(\tilde{x}^{\prime})$. The lifted backprojection $G_{i}^{a}(x)$ for a state $x$ and action $a$ then sums $g_{i}^{\tilde{a}}(\tilde{x})$ for each possible propositional assignment $\tilde{x}$ and $\tilde{a}$ and weights the term with the counts given by the state $x$.
@@
Approximate Foreplan precomputes all backprojections and then builds the following linear program [16]:
$\begin{array}[]{ll@{}}\text{Variables:}&w_{1},\dots,w_{n}\ ;\\ \text{Minimize:}&\displaystyle\sum_{i=1}^{n}\alpha_{i}w_{i}\ ;\\ \text{Subject to:}&\forall a\in A:\\ &0\geq\displaystyle\max_{x}\left\{R(x)+\displaystyle\sum_{i=1}^{n}w_{i}\left(\gamma G_{i}^{a}(x)-h_{i}(x)\right)\right\}.\\ \end{array}$ (5)
The $\alpha_{i}$’s are effectively coefficients for a linear combination over the $w_{i}$, stating how important the minimization of each $w_{i}$ is [16, 8]. The maximum operator is not part of the definition of linear programs and is removed in an operation similar to variable elimination (VE) [35]. In Appendix A.4, we provide an example for the removal procedure.
For the runtime analysis, we introduce the cost network briefly: The cost network for a constraint has a vertex for each appearing variable and there is an edge in the cost network between two vertices if the corresponding variables appear together in the same function.
Theorem 17.
Proof.
Approximate Foreplan has to solve the linear program in Equation 5. The number of variables and constraints in the linear program is linear in the action space and exponential in the induced width of each cost network [11]. Because Approximate Foreplan does not iterate over the whole state space, but treats each clique independently in the maximum operator, the effective state space is no longer exponential in $c$, but polynomial, and the action space is bound by the state space. ∎
@@
Moreover, the same approximation guarantee as for ALP holds for Approximate Foreplan. ALP provides the best approximation of the optimal value function in a weighted $\mathcal{L}_{1}$ sense, where the weights in the $\mathcal{L}_{1}$ norm are the state relevance weights $\alpha$ [16, 8]. To show that the results can be carried over, we show that Approximate Foreplan and ALP are equivalent on rfMDPs:
Theorem 18.
@@
With Approximate Foreplan, we reduced the runtime further from exponential in $c$ to polynomial in $c$. In the next section, we extend (Approximate) Foreplan to identify how many objects an agent should act on to achieve a certain task given restrictions.
7 Conditional Action Queries
Foreplan can be used, e.g., by a mayor to find the optimal number of persons to impose a travel ban on. But sometimes, the mayor could be interested in how she can achieve that the probability of at least half of the town’s population being healthy is at least $p$. Therefore, let us first define this query type formally:
Definition 19 (Conditional Action Query).
A conditional action query for a state $x$ in an rfMDP is a threshold $t\in\mathbb{R}\cup\{-\infty\}$ together with a restriction query $P(\cdot\mid x,a)\geq p$ with a probability threshold $p$. The answer to a conditional action query is the action $a$ the agent has to perform to get an expected reward of at least $t$ in the next state and fulfilling the restriction query.
To answer a conditional action query, we have to compute the maximum over all actions for the one-step lookahead
$Q_{a}(s)=R(s)+\gamma\sum_{s^{\prime}}P(s^{\prime}\mid s,a)\cdot V(s^{\prime})$ (6)
or, for Approximate Foreplan,
$Q_{a}(x)=R(x)+\gamma\sum_{i}w_{i}g_{i}^{a}(x).$ (7)
We can use the techniques from (Approximate) Foreplan to calculate $Q_{a}(x)$ efficiently. Then, we know for every action the expected reward in state $x$ and keep the ones where the expected reward is at least $t$. We further filter the actions by checking the restriction query $P(\cdot\mid x,a)\geq p$ with a call to Lifted Variable Elimination (LVE) [26], providing the state $x$ and action $a$ as evidence.
Theorem 20.
Given that the (I) results of Foreplan are available, (II) parfactor model of the rfMDP is liftable, and (III) evidence for state and action is liftable, we can answer conditional action queries in time (i) exponential in $c$ and $w$, and (ii) polynomial in the number of objects and action space. With approximation, we are not exponential, but polynomial in $c$ and in the number of basis functions.
@@ 7 Conditional Action Queries
Equation 6 has to be computed for all possible actions. Thus, it requires iteration over the whole action and state space. For the approximation, Equation 7 again has to be computed for all possible actions. However, the state and action spaces are now only polynomial in the number of basis functions and not exponential in $c$. The number of backprojections is the same as of the basis functions. Last, LVE has to be run to check the restriction query for each action. LVE runs in time polynomial in the number of objects if the parfactor model is liftable and evidence for state and action is liftable. ∎
Theorem 20 enables iterative planning. We execute Foreplan once and reuse the results to answer conditional action queries at each timestep, leaving us with the time complexity stated in Theorem 20. Furthermore, one could adjust Foreplan to take the conditional action query into account when planning. This way, Foreplan readily returns the answer to the query for each state. Moreover, most typically, $c$ and $w$ are much lower than the number of objects $n$ and independent of $n$, that is, almost a constant. Furthermore, the runtime in $n$ is of much more interest:
Corollary 21.
The runtime for answering conditional action queries is polynomial in the number of objects with $c$ and $w$ as constants and if the parfactor model and the evidence are both liftable.
This result sets Foreplan greatly apart from traditional approaches requiring time exponential in $n$. Further, it gives rise to an iterative planning framework: At each time, the mayor performs an action keeping more than half of population healthy with a high probability. The decision can be made using Foreplan to evaluate the possible actions. In the next section, we evaluate Foreplan empirically.
8 Empirical Evaluation
@@
(Approximate) Foreplan runs in time polynomial in the number of objects, but with other terms are unavoidably exponential. In contrast to current approaches, the exponential terms of both Foreplane variants depend only on the structure of the rfMDP and not on the number of objects. To underline our theoretical results, we evaluate (Approximate) Foreplan against ALP and an implementation of symbolic value iteration using extended algebraic decision diagrams (XADDs) [18, 32] for the epidemic example introduced in Example 3. We use Python 3.12 and HiGHS for solving the linear programs [17]. We run all implementations on a 13th Gen Intel(R) Core(TM) i5-1345U with 1.60 GHz and 16 GB of RAM.
@@
Figure 2 shows the runtime for the epidemic example for (Approximate) Foreplan, ALP and XADD Symbolic Value Iteration for up to 22 persons and Figure 3 in Appendix A.7 shows the runtimes for up to 164 persons, both with a time limit of two hours. XADD Symbolic Value Iteration exceeds the time limit after eight persons, ALP after 15 persons. Foreplan runs after 21 persons and Approximate Foreplan after 164 persons out of memory. For eight persons, Foreplan is $100$ times faster than XADD Symbolic Value Iteration. Approximate Foreplan is even $10.000$ times faster, which are four orders of magnitude. For 15 persons, Foreplan is more than six times faster than ALP and Approximate Foreplan is more than $8.576$ times faster than ALP. For 21 persons, Approximate Foreplan is more than $4.963$ times faster than Foreplan. Thus, when using a symbolic solver, we can only solve the epidemic example for up to eight persons. With a factored and approximate approach, we can go up to 15 persons. In contrast, when using Foreplan, we can solve the epidemic example even for 21 persons and with Approximate Foreplan we can go further to 164 persons, which is ten times more than what ALP can solve. While Foreplan appears to be rising exponentially in the number of persons in Figure 2, we note that by Theorem 15 it is polynomial in the number of persons, which can be seen in Figure 3 in Appendix A.7. Overall, (Approximate) Foreplan achieves a speedup of several orders of magnitude and is able to compute policies for significantly more persons within the same time and memory limits.
@@
Figure 2: Runtime of (Approximate) Foreplan, ALP and XADD Symbolic Value Iteration on the epidemic example for up to 22 persons with a time limit of two hours.
9 Conclusion
@@
Propositional planning approaches struggle with having numerous indistinct objects and actions on subsets of those objects. While first-order MDPs can cope with numerous objects, they still need to represent actions for each subset individually, resulting in exponentially many actions. In this paper, we present Foreplan, a relational forward planner solving the exponential explosion by lifting the objects: Foreplan groups indistinguishable objects by a representative and carries out calculations on a representative-level. Afterwards, the result is projected to the whole group. Using histograms and focusing only on the number of objects an action is applied to, we effectively reduce the action space from exponential to polynomial in the number of objects. Foreplan allows for answering conditional action queries, enabling iterative planning with constraints. Such queries can be solved in time polynomial in the number of objects. Moreover, Foreplan opens new doors for mechanism design for incentivizing the concurrent actions in the returned policy.
@@
In future work, we aim to develop a hybrid approach combining Foreplan and Golog [23]. With the forward search in Foreplan, we can identify states reachable from the initial state while the backwards search in Golog computes the exact optimal policy. Furthermore, the techniques from Foreplan can be transferred to first-order partially observable MDPs [34].
{ack}
@@ Appendix A Omitted Details
A.1 Examples for the Relational Cost Graph

### v2

@@ 5. Complexity Analysis of Foreplan
Note that $c$ and $w$ are independent of the domain sizes and determined only by the structure of the relational cost graph, and thus in general small. Also, $c$ is bounded by the number of PRVs in our parfactor model, as one parfactor is sufficient per PRV. Theorem 5 overapproximates the size of the state representation, as not all cliques have the same size and not both, $c$ and $w$ are large at the same time. Building on the size of the state representation, we give the complexity of the state and action space: {theorem} The state and action spaces are both polynomial in the number of objects and exponential in $c$ and $w$.
@@
We need to iterate over all possible instantiations of the state representation. For each clique (resp. CRV), the number of possible instantiations is polynomial in the number of objects and exponential in $w$. The joint state requires one instantiation per clique, resulting in the number $c$ of CRVs as exponent. The size of the action space is bounded by the size of the state space, as the action has to specify a (subset of a) state. ∎
Since Foreplan uses a linear program to compute the value function, we analyze the complexity of solving the linear program Foreplan builds. Linear programs can be solved in polynomial time w.r.t the variables and constraints 63499. Let us therefore take a closer look at the number of constraints and variables Foreplan generates:
{theorem}
@@
Plugging Theorem 5 in Theorem 5 leads to: {theorem} The runtime of Foreplan is polynomial in the number of objects and exponential in $c$ and $w$.
Therefore, with Foreplan, we have introduced an exact planner with a runtime polynomial instead of exponential w.r.t domain sizes for concurrent actions. The other parameters Foreplan is exponential in are mostly small and fixed. Thus, we have already achieved an exponential speed up.
Our main contribution Foreplan greatly advances the state of the art. However, as it is an exact algorithm, the value function is still not factorized. In case the runtime is of utter importance, we can also approxiamte the value function, following the idea of approx_factored_mdps, which has also been picked up by, e.g., 10.5555/3020336.3020398. By approximating the value function, we can prevent iterating over the joint state space leading to a blazingly fast, but approximate, version of Foreplan.
6. Foreplan: Even Faster by Approximation
While Foreplan runs in time polynomial in the number of objects, the runtime still depends exponentially on $c$. In this section, we present an approximation technique inspired by the Approximate Linear Programming (ALP) approx_factored_mdps approach to prevent iterating the whole state space. Our approach follows the same idea as ALP and ALP for first-order MDPs 10.5555/3020336.3020398. We first describe the approximation idea and then how Foreplan uses the approximation for our new representation and for concurrent actions. Last, we give bounds on the runtime and on the approximation quality. We call the approximate version Approximate Foreplan. Since Approximate Foreplan is an extension of our main contribution Foreplan, we refer to the Appendix for technical details.
Foreplan needs to iterate over the whole state space, because the value function maps from a state to the value of that state. We approximate the value function by a set $h_{i}$ of basis functions, whose scope is a subset of $\mathbf{S}$, $V\approx\sum_{i}w_{i}h_{i}$, where the goal is to find the most suitable weights $w_{i}$ approx_factored_mdps. Approximate Foreplan also needs the value of all possible next states in terms of the same approximation. Therefore, Approximate Foreplan makes use of the backprojections $g_{i}^{a}$ of the basis functions $h_{i}$ approx_factored_mdps, stating the influence of $x$ on the next state:
@@ 6. Foreplan: Even Faster by Approximation
To compute the basis functions and backprojections lifted, we parameterize the basis functions analogue to the reward functions and calculate them lifted in the same way. The basis functions should capture the important dynamics in the model koller1999computing, most importantly the rewards 10.5555/3020336.3020398. As proposed by koller1999computing and picked up in 10.5555/3020336.3020398, we also use basis functions for capturing the rewards alongside a constant basis function:
[…]
The backprojections are computed lifted as defined next:
[…]
@@
Approximate Foreplan precomputes all backprojections and then builds the following linear program approx_factored_mdps, whose result is the approximated value function and constitutes the result of Approximate Foreplan:
[…]
The $\alpha_{i}$’s are effectively coefficients for a linear combination over the $w_{i}$, stating how important the minimization of each $w_{i}$ is approx_factored_mdps; doi:10.1287/opre.51.6.850.24925. The maximum operator is not part of the definition of linear programs and is removed in an operation similar to variable elimination (VE) ve. In Appendix A.6, we provide an example for the removal procedure. Moreover, in Appendix B, we show a walkthrough of Approximate Foreplan on the epidemic example.
For the runtime analysis, we introduce the cost network briefly: The cost network for a constraint has a vertex for each appearing variable and there is an edge in the cost network between two vertices if the corresponding variables appear together in the same reward or basis function, or backprojection. For the complexity analysis, we use the structural graph parameter induced width, which provides an upper bound for the largest intermediate result dechter1999bucket:
{theorem}
Approximate Foreplan runs in time polynomial in the number of objects, polynomial in $c$ and exponential in the induced width of each cost network, when $w$ is bounded.
Proof Sketch.
The full proof is in Appendix A.7. Approximate Foreplan has to precompute the backprojections and solve the linear program in Equation 5. The first involves constantly many iterations of the state and action spaces. The second is linear in the action space and exponential in the induced width of each cost network approx_factored_mdps; dechter1999bucket. Because Approximate Foreplan does not iterate over the whole state space, but treats each clique independently in the maximum operator, the effective state space is no longer exponential in $c$, but polynomial, and the action space is bound by the state space. ∎
@@
Moreover, we can show that Approximate Foreplan and ALP compute the same solutions: {theorem} Given an rfMDP $R$, Approximate Foreplan and ALP are equivalent on $R$ and the grounded version of $R$, respectively.
@@
With ALP and Approximate Foreplan being equivalent, we can transfer the approximation guarantee for ALP to Approximate Foreplan:
[…]
Approximate Foreplan provides the best approximation of the optimal value function in a weighted $\mathcal{L}_{1}$ sense, where the weights in the $\mathcal{L}_{1}$ norm are the state relevance weights $\alpha$.
@@
Since the claim holds for ALP approx_factored_mdps; doi:10.1287/opre.51.6.850.24925, the proof follows by the equivalence of ALP and Approximate Foreplan. ∎
With Approximate Foreplan, we reduce the runtime further from exponential in $c$ to polynomial in $c$. In the next section, we evaluate the runtimes of Foreplan and Approximate Foreplan empirically.
7. Empirical Evaluation
@@
(Approximate) Foreplan runs in time polynomial in the number of objects, but other terms are unavoidably exponential. In contrast to current approaches, the exponential terms of both Foreplan variants depend only on the structure of the rfMDP and not on the number of objects. To underline our theoretical results, we evaluate (Approximate) Foreplan against ALP and an implementation of symbolic value iteration (VI) using extended algebraic decision diagrams (XADDs) pmlr-v162-jeong22a; taitler2022pyrddlgym for the epidemic example introduced in Example 3 as well as for the BoxWorld boutilier2001symbolic and a fully-connected SysAdmin approx_factored_mdps example, with the latter two extended to concurrent actions. We also assess the quality of the policy given by Approximate Foreplan. We use Python 3.12 and HiGHS for solving the linear programs highs. We run all implementations on an AMD EPYC 7452 32-Core Processor with 504 GB of RAM.
@@
Figure 4 shows the runtime for the epidemic example for (Approximate) Foreplan, ALP and XADD Symbolic VI for up to 191 persons with a time limit of two hours. XADD Symbolic VI exceeds the time limit after ten persons, ALP after 15 persons. Foreplan times out after 20 persons and Approximate Foreplan after 191 persons. For ten persons, Foreplan is $91$ times faster than XADD Symbolic VI. Approximate Foreplan is even $14.000$ times faster, which are four orders of magnitude. For 15 persons, Foreplan is more than four times faster than ALP and Approximate Foreplan is more than $4.217$ times faster than ALP. For 20 persons, Approximate Foreplan is more than $4.077$ times faster than Foreplan. Thus, when using a symbolic solver, we can only solve the epidemic example for up to ten persons. With a factored and approximate approach, we can go up to 15 persons. In contrast, when using Foreplan, we can solve the epidemic example even for 20 persons and with Approximate Foreplan we can go further to 191 persons, which is ten times more than what ALP can solve.
We also run the four algorithms on a fully connected SysAdmin example with a timeout of two hours. XADD Symbolic VI times out after eleven computers and ALP after 12. With Foreplan, we can go up to 64 computers and with Approximate Foreplan even up to 94 computers. At eleven computers, Foreplan is more than $44.413$ times faster than XADD Symbolic VI and Approximate Foreplan is even more than $69.944$ times faster. At 12 computers, Foreplan is more than $8.359$ times faster than ALP and Approximate Foreplan even more than $16.173$ times. For 64 computers, Approximate Foreplan is more than $696$ times faster than Foreplan. In Appendix A.10, we present the runtime figure.
For the BoxWorld example, we use three cities and set a time limit of 15 hours. XADD Symbolic VI times out after two boxes and trains. ALP and Foreplan time out after nine boxes and trains. Approximate Foreplan manages to go up to 30 boxes and trains. At two boxes and trains, Foreplan is more than $83.227$ times faster than XADD Symbolic VI and Approximate Foreplan even more than $137.502$ times. At nine boxes and trains, Foreplan is more than three times faster than ALP and Approximate Foreplan is even more than $33.136$ times faster than ALP. In Appendix A.10, we present the runtime figure.
For assessing the quality of the policy given by Approximate Foreplan, we calculate the ratio of wrong actions over the total number of actions in every ground state. For ten persons, ALP and Approximate Foreplan both deviate in $1.2$ % of all actions. For two to ten persons, the deviation was $2.98$ % at most, with reducing errors when increasing the number of individuals. In all cases, ALP and Approximate Foreplan have an identical error. For SysAdmin, ALP and Approximate Foreplan both return the optimal policy in the tests going up to nine computers. For BoxWorld, we refer to Appendix A.10.
Overall, (Approximate) Foreplan achieves a speedup of several orders of magnitude and is able to compute policies for significantly more objects within the same time and memory limits. Moreover, the policy given by Approximate Foreplan is for more than $97$ % of all actions in the epidemic example correct and for all actions in the SysAdmin example.
@@
Figure 4. Runtime of (Approximate) Foreplan, ALP and XADD Symbolic VI on the epidemic example for up to 191 persons with a time limit of two hours.
[…]
Graphical runtime comparison of Foreplan, Approximate Foreplan, Approximate Linear Programming and XADD Symbolic VI.
8. Conclusion
@@
Propositional planning approaches struggle with having numerous indistinct objects and actions on subsets of those objects. While first-order MDPs can cope with numerous objects, current approaches for solving planning in them neglect concurrent actions and still need to represent actions for each subset individually, resulting in exponentially many actions. In this paper, we present Foreplan, a relational forward planner solving the exponential explosion by lifting the objects: Foreplan groups indistinguishable objects by a representative and carries out calculations on a representative-level. Afterwards, the result is projected to the whole group. Using histograms and focusing only on the number of objects an action is applied to, we effectively reduce the action space from exponential to polynomial in the number of objects.
@@
In future work, we aim to develop a hybrid approach combining Foreplan and Golog levesque1997golog. With the forward search in Foreplan, we can identify states reachable from the initial state while the backwards search in Golog computes the exact optimal policy. Furthermore, the techniques from Foreplan can be transferred to first-order partially observable MDPs williams2005factored.
{acks}
@@ Appendix A Omitted Details
A.1. Examples for Hyperoperations in Relational Factored MDPs
Definition 3 requires the rewards to be represented as a sum of local reward functions. In fact, this requirement can be lifted to a broader one: Any operation $\oplus$ can run over a collection of local reward functions, as long as this operation has a hyperoperation $\bigoplus$, which performs $\oplus$ repeatedly. The most basic example is addition ($\oplus=+$) and multiplication as the corresponding hyperoperation ($\bigoplus=\cdot$). Moreover, these operations can be linked together in any way: We compute each operation on a collection of local reward functions lifted and then compute the overall result following the linking.
Let us first define some local reward functions and the reward function:
[…]
Assume we have PRVs $A(M)$ and $B(M)$, both with Boolean domain. We define the local reward functions $r_{1}(A(M))$ with $r_{1}(false)=2$ and $r_{1}(true)=5$. We further define $r_{2}(B(M))$ with $r_{2}(false)=-1$ and $r_{2}(true)=1$. We define the reward function as $\sum_{m\in M}r_{1}(m)\cdot\prod_{m\in M}r_{2}(m)$
The lifted computation of the reward function then calculates each basic term, i.e., $\sum_{m\in M}r_{1}(m)$ and $\prod_{m\in M}r_{2}(m)$, separately and combines the results using the linking operations, which is in this example the multiplication:
[…]
Assume $\left\lvert\mathcal{D}(M)\right\rvert=5$ and $A(M)$ is true for two of them and $B(M)$ is true for three of them. The lifted computation of the first term is then:
$\sum_{m\in M}r_{1}(m)=2\cdot r_{1}(true)+3\cdot r_{2}(false),$ (6)
and for the second term:
$\prod_{m\in M}r_{2}(M)=\left(r_{2}(true)\right)^{3}\cdot\left(r_{2}(false)\right)^{2}.$ (7)
The overall result is thus:
$\left(2\cdot r_{1}(true)+3\cdot r_{2}(false)\right)\cdot\left(\left(r_{2}(true)\right)^{3}\cdot\left(r_{2}(false)\right)^{2}\right).$ (8)
The hyperoperation is needed to reuse the result of one representative, lifting it to the group.
A.2. Formal Definition of Relational Cost Graph
To give a formal representation of the relational cost graph, we first define an undirected graph:
[…]
A graph $G=(V,E)$ is tuple of a set $V$ of vertices and $E\subseteq V\times V$ of edges. We require $E$ to be symmetric, i.e., $(a,b)\in E\implies(b,a)\in E$.
We further need some shorthand syntax to extract the logvars a PRV is defined over and to extract the PRVs a parfactor or local reward function is defined over:
[…]
Given a PRV $B$ defined over logvars $L_{1},\dots,L_{n}$, the notation $args(B)$ returns the logvars $L_{1},\dots,L_{n}$, i.e.,
$args(B)\coloneqq\{L_{1},\dots,L_{n}\}.$ (9)
We analogously define argsargs to return the arguments of a parfactor or local reward function: Given a parfactor $\phi$ defined over PRVs $B_{1},\dots B_{m}$, $args(\phi)$ returns those PRVs:
$args(\phi)\coloneqq\{B_{1},\dots,B_{m}\}.$ (10)
And given a local reward function $r$ defined over PRVs $A_{1},\dots,A_{k}$, $args(r)$ returns those PRVs:
$args(r)\coloneqq\{A_{1},\dots,A_{k}\}.$ (11)
Using these shorthand notations, we can give a formal definition of the relational cost graph:
[…]
Given an rfMDP with a set $\mathbf{R}$ of parameterized local reward functions, a set $\mathbf{\Phi}$ of factors, and a set $\mathbf{X}$ of PRVs, the relational cost graph is a graph $rcg=(V,E)$ with $V\coloneqq\mathbf{X}$ and the set $E$ of edges given by
$E\coloneqq\{(A,B)$ $\in\mathbf{X}\times\mathbf{X}\mid\left(args(A)\cap args(B)\neq\emptyset\right)$ (12)
$\land\left(\exists\phi\in\mathbf{\Phi}:a\in args(\phi)\land b\in args(\phi)\right)$ (13)
$\land\left(\exists r\in\mathbf{R}:a\in args(r)\land b\in args(r)\right)\}.$ (14)
A.3. Examples for the Relational Cost Graph